A 506 g mass oscillates with an amplitude of 11.0 cm on a spring whose spring constant is 20.0 n/m. Determine the period. Determine the maximum speed.
Answer #1

Given,
Mass, m = 506 g = 0.506 kg
Amplitude, A = 11 cm = 0.11 m
Spring constant, k = 20 N/m

Time period,
T = 2\pi\sqrt(k/m)
T = 2\pi\sqrt(20/0.506)
T = 39.48 s

Maximum velocity,
vmax = A\sqrt(k/m)
vmax = 0.11\sqrt(20/0.506)
vmax = 0.69 m/s

_______________________

Please give your valuable feedback

Thank you